OpenStudy (anonymous):

Choose the slope-intercept equation of the line that passes through the point (-2, 2) and is parallel to y = 4x + 7. y = 4x + 10 y = one fourthx + 5 halves y = -4x - 6 y = -one fourthx + 3 halves

OpenStudy (anonymous):

i need help

OpenStudy (anonymous):

Parallel means they'll have the same slope the slope you need in your new equation needs to be 4 to find an equation with that point in it use this equation \[y-y_1=m(x-x_1)\] where m=slope and \[(x_1,y_1)=(-2,2)\]

OpenStudy (anonymous):

Can you plug those in?

OpenStudy (anonymous):

Are you there? I'm trying to help you...

OpenStudy (anonymous):

sorry im back

OpenStudy (anonymous):

Okay. substitute those in the given equation and I can help you finish up the problem!

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so (-1,1)?

OpenStudy (anonymous):

What?

OpenStudy (anonymous):

I only asked you to put the information we new into the equation and you gave me a point.

OpenStudy (anonymous):

\[y-(2)=(4)(x-(-2))=\]\[y-2=4(x+2)\] Use distributive property \[y-2=4x+8\] Get y itself by adding 2 on both sides \[y=4x+10\]

OpenStudy (anonymous):

ok